Chapter 3 Problems Saved Help Save & Exit Submit Check my work 9 You are a manager in charge of monitoring cash flow at a major publisher. Paper books comprise 40 percent of your revenues, which grow about 2 percent annually. You recently received a preliminary report that suggests the growth rate in ebook reading has leveled off, and that the cross-price elasticity of demand between paper books and ebooks is -0.3. In 2016, your company earned about $600 million from sales of ebooks and about $400 million from sales of paper books. 10 points If the own price elasticity of demand for paper books is -2, how will a 4 percent decrease in the price of paper books affect your overall revenues from both paper books and ebooks sales? eBook Print References Instruction: Enter your response rounded to one decimal place. Your overall revenues will change by $million Prev 9 of 9 Next
Answer #1

Own price elasticity = % Change in quantity demanded of paper books / % Change in price of paper books

-2 = % Change in quantity demanded of paper books / (-4%)

% Change in quantity demanded of paper books = (-2) x (-4)% = 8% (increase)

Let initial price and quantity of paper books be P0 and Q0.

Initial revenue = P0 x Q0 = $400 million

New revenue = (0.96 x P0) x (1.08 x Q0) = (0.96 x 1.08) x (P0 x Q0) = 1.0368 x $400 million = $414.72 million

Increase in revenue = $(414.72 - 400) million = $14.72 million............(1)

Also,

Cross-price elasticity = % Change in quantity demanded of e-books / % Change in price of paper books

-0.3 = % Change in quantity demanded of e-books / (-4%)

% Change in quantity demanded of e-books = (-0.3) x (-4)% = 1.2% (increase)

Let initial price of paper books and quantity of e-books be P0 and Q0.

Initial revenue = P0 x Q0 = $600 million

New revenue = (0.96 x P0) x (1.012 x Q0) = (0.96 x 1.012) x (P0 x Q0) = 0.97152 x $600 million = $582.912 million

Decrease in revenue = $(600 - 582.912) million = $17.088 million............(2)

Overall change in revenue = (1) - (2) = $(14.72 - 17.088) million = -$3.088 million = -$3.1 million